OpenStudy (anonymous):

9. In his free time, Gary spends 7 hours per week on the Internet and 14 hours per week playing video games. If Gary has five hours of free time per day, what percent of his free time is spent on the Internet and playing video games?

OpenStudy (anonymous):

are you in connections academy?

OpenStudy (mathstudent55):

7 hours + 14 hours = 21 hours of free time Internet: 7/21 = 1/3 = 33.33... % Games: 14/21 = 2/3 = 66.66... %